solve for all values of x that satisfy the equation x/x+3=5/x+7

Answer by madhan_math(34)

x/x+3=5/x+7
solution:
Taking x as the L.C.M. of both side then denominator x will be cancelled.
Now we have x+3x=5+7x.
This implies -3x=5.
Therefore x has the value x=-5/3.